# Ax-Prover: A Deep Reasoning Agentic Framework for Theorem Proving in Mathematics and Quantum Physics

### v2

@@ 7 Conclusions
Evaluations show that Ax-Prover ranks third on the challenging PutnamBench and outperforms baselines on the public NuminaMath-LEAN dataset as well as on AbstractAlgebra and QuantumTheorems, two new datasets we introduce that focus on research-level mathematics and physics. These benchmarks not only provide new testbeds for cross-domain reasoning in future agents but also represent a crucial milestone in evaluating reasoning models in any scientific discipline grounded in mathematics.
@@
Furthermore, we showed in our case study on cryptography that Ax-Prover is not only able to prove theorems autonomously, but it can also engage in fruitful collaboration with human researchers. Working alongside it, the mathematician used it as a partner for structuring arguments, validating intermediate lemmas, and diagnosing proof failures. This interaction demonstrates how Ax-Prover can adapt to expert guidance, accelerate verification workflows, and even surface errors in the informal reasoning.
Looking ahead, we plan to enhance Ax-Prover by introducing parallelized agents, enabling the framework to explore multiple proof paths simultaneously. This will increase its creativity and success rate in formalizing complex proofs. We also plan to integrate a long-term memory module to store information from past proofs and human interactions. This capability will allow Ax-Prover to participate not only in standalone problems but also in extended, collaborative research projects requiring sustained expert guidance. These developments will advance us towards our broader goal of verifiable scientific artificial intelligence, where AI systems contribute to scientific discovery through formally validated reasoning.
@@ B.1.1 Dataset Generation
We used a basic pipeline to build the AbstractAlgebra dataset. First, we extracted all raw text from PDFs of Abstract Algebra by Dummit and Foote [23] and Abstract Algebra: Theory and Applications by Judsen [29] using Mistral’s API. We then processed the raw text by using Claude-Sonnet-3.7 [3] to extract a list of natural language mathematical statements, which include exercises, derivations, lemmas, propositions, and theorems within the text. Next, we used a Claude-Sonnet-3.7 agent to formalize each natural language statement in Lean. To ground the formalization in Mathlib and prevent the agent from reinventing definitions, we passed the agent a Lean file at the start of the process containing relevant definitions for that section, e.g., dihedral groups, roots of unity, or the field extension $\mathbb{Q}(\sqrt{2})$. The agent could reference these definitions and was required to add each formalized statement directly to this file, but was explicitly prohibited from introducing new definitions. The agent generated the top 3 formal statements in Lean for each natural language statement and refined each attempt up to 3 times with feedback from the Lean compiler. We then built the dataset by retaining only those pairs of natural language and formal language statements that corresponded to exercises from the source texts.
@@ B.1.2 Example
This is an example of a theorem statement in the AbstractAlgebra dataset, formalized from Exercise 3 in Chapter 1.2 of Dummit and Foote [23].
@@
/-- Use the generators and relations to show that every element of D_n not a power of r has order 2. -/
@@ B.2.1 Dataset Generation
The dataset was generated through an iterative human-in-the-loop process combining automated proof synthesis with expert curation. The initial set of 134 theorems were manually extracted from [43]. An automated coding agent (Claude Opus [6]) first generated formal statements and proof attempts for the theorems, producing both complete proofs and partial derivations. A quantum physics expert then reviewed each statement and proof, identifying gaps, correcting errors, and standardizing operator definitions to ensure that each question was well formed and solvable. The final dataset replaces these proofs with sorry statements.
@@ Appendix C Detected Autoformalization Error
As noted in Section 5.2, 19.7% of Numina’s problems were generated using autoformalization models. While these pipelines enable large-scale dataset construction, they occasionally produce ill-posed theorems that cannot be satisfied in Lean.
@@ Appendix E Tactics
Tactic Ax-agent DeepSeek Kimina
@@
Table 5: Tactics used by Ax-agent, DeepSeek, and Kimina. An ”X” indicates the model uses the tactic.
@@ Appendix F Case Study: Verifying math in classical cryptographic papers
As a concrete example, we focus on the recent (May 2024) cryptographic paper A New Algorithm for Computing Branch Number of Non-Singular Matrices over Finite Fields from arXiv [40]. This work introduces a novel algorithm for computing the branch number – a fundamental metric used to assess the strength of block ciphers such as AES [42], PRINCE [12], and Grøst [24].
@@
For cryptographers, this makes a practical difference: it enables fast evaluation of candidate matrices when designing new lightweight or high-performance ciphers. The authors demonstrate in Theorem 4 [40] that their algorithm achieves significant complexity gains over the naive $O(n^{2}q^{n})$ approach for finite fields of order $q\geq 4$ and square matrices of order $n\geq 4$.
@@ F.4 Correctness, Verification, and Takeaways
The outcome of our work is a formally verified certificate ensuring the correctness of Theorem 1. During the course of constructing this certificate, we identified a critical error in the original proof. Specifically, the issue arises in Step 2 of the proof:
@@
This matters for two reasons:
  1. 1.
    Logical correctness: Reasoning about the empty set is problematic (all statements are vacuously true) which can lead to unsound conclusions. For example, let
    $S=\{\,x\in\mathbb{Z}\mid x=3\ \text{and}\ x\ \text{is even}\,\}.$
    Take $y\in S$, then $y=3$ and $y$ is even, so this implies that 3 is even.
  2. 2.
    Software implementation: Computing the minimum of an empty set is undefined in standard programming environments and would trigger a runtime error if translated directly into code.
Our formal verification system flagged these issues because it could not establish the truth of the corresponding statements, revealing logical gaps in the proof. Nevertheless, the authors’ final result remains correct despite the critical error in their proof.
Appendix G Case Study 2: Quantum Cryptography
Quantum cryptography seeks security guarantees that are statistical; derived from the laws of physics and information theory, rather than computational. An important example is quantum key distribution (QKD), a key–establishment protocol in which two parties certify secrecy by testing quantum correlations rather than assuming limits on an adversary’s computing power. Because these guarantees rest on first principles in linear algebra, probability, and quantum mechanics, the field is an especially natural fit for automated theorem proving: formal proofs can turn widely cited derivations into reusable components that compose into end‑to‑end security arguments.
Our second use case focuses on the Lo–Chau security framework [37], reproduced in this work. That paper laid the foundations of modern QKD by reducing security to verifiable statements about entanglement and then to classical probabilistic reasoning, and it informed subsequent analyses such as the Shor–Preskill proof of BB84 [52]. Within that framework, a key step is to convert a physical test, high fidelity to $R$ EPR pairs, into an information‑theoretic guarantee of low entropy and therefore limited eavesdropper information.
Using Ax‑Prover in interactive mode, we proved in Lean the entropy bound that implements this conversion, Lo–Chau’s Lemma 1 (High fidelity implies low entropy), and packaged it as a library lemma for downstream use (see Section G). This result illustrates how a tool‑based prover can assist domain experts in translating physics‑style arguments into machine‑checked mathematics, strengthening the statistical‑security foundations on which modern quantum cryptography is built.
In what follows, we cite and explain in detail the lemma and corresponding lean certificate.
G.1 Quantum Key Distribution Lemma
Let $\rho$ be a density matrix on a $2^{2R}$‑dimensional Hilbert space. If its fidelity with the ideal RR‑singlet state satisfies $\langle R\text{~singlets}\mid\rho\mid R\text{~singlets}\rangle>1-\delta$ with $0<\delta\ll 1$, then
@@ G.1 Quantum Key Distribution Lemma
This is the form reproduced in this work from the Lo–Chau reprint [37]. Here, however, instead of the statement $\delta\ll 1$ we used: $\delta\leq 1-\frac{1}{2^{2R}-1})$
The fidelity condition implies the largest eigenvalue of $\rho$ is at least $1-\delta$. Since von Neumann entropy is Schur‑concave, the maximal entropy under this constraint is achieved by the extremal spectrum $\bigl(1-\delta,\tfrac{\delta}{2^{2R}-1},\ldots,\tfrac{\delta}{2^{2R}-1}\bigr)$, which yields the stated bound. Our Lean certificate for this result follows this reduction and checks the necessary concavity and normalization facts. This lean certificate can be made available upon request.
In the Lo–Chau security reduction, the lemma turns “almost‑perfect EPR pairs” into a quantitative entropy bound that, together with standard information‑theoretic tools, limits an eavesdropper’s knowledge of the final key. Making this step machine‑checked enables principled composition with other verified components in formal QKD security proofs. Thus, Ax-Prover bridges formal reasoning and quantitative quantum information theory: results such as the Lo–Chau entropy bound no longer have to be taken as a black box, but instead become certified components, ready for use in end-to-end formal verification of QKD security proofs.

### v3

@@ 6 Conclusions
Evaluations show that Ax-Prover is the top performing open-source model and third overall on PutnamBench, using a much lower computation budget than top-performers, and outperforms baselines on the public NuminaMath-LEAN dataset as well as on AbstractAlgebra and QuantumTheorems, two new datasets we introduce that focus on research-level mathematics and physics. These benchmarks not only provide new testbeds for cross-domain reasoning in future agents but also represent a crucial milestone in evaluating reasoning models in any scientific discipline grounded in mathematics.
@@
Furthermore, we showed in our case studies that Ax-Prover is not only able to prove theorems autonomously, but also to engage in fruitful collaboration with researchers. Working alongside it, they used it as a partner for structuring arguments, validating lemmas, and diagnosing proof failures. This interaction demonstrates how Ax-Prover can adapt to expert guidance, accelerate verification workflows, and even surface errors in the informal reasoning.
Looking ahead, we plan to enhance Ax-Prover by introducing parallelized agents, enabling the framework to explore multiple proof paths simultaneously. This will increase its creativity and success rate in formalizing complex proofs. We also plan to integrate a long-term memory module to store information from past proofs and human interactions. This capability will allow Ax-Prover to participate not only in standalone problems but also in extended, collaborative research projects. These developments will advance us towards our broader goal of verifiable scientific artificial intelligence, where AI systems contribute to scientific discovery through formally validated reasoning.
Acknowledgments
We thank Austin Letson for reviewing the paper, suggesting that we evaluate Ax-Prover on PutnamBench, and assisting with Lean-related tasks, including the use of lean4checker to verify our results. We thank Avinash Kumar for directing us to the NuminaMath-LEAN dataset. Finally, we acknowledge Leopoldo Sarra for working closely with us during the early stages of the project.
Appendix
@@ B.1.1 Dataset Generation
We used a basic pipeline to build the AbstractAlgebra dataset. First, we extracted all raw text from PDFs of exercises from Abstract Algebra by Dummit and Foote [dummit2004abstract] using Mistral’s API. We then processed the raw text by using Claude Sonnet 3.7 [anthropic2025claude37] to extract a list of natural language mathematical statements, which include exercises, derivations, lemmas, propositions, and theorems within the text. Next, we used a Claude Sonnet 3.7 agent to formalize each natural language statement in Lean. To ground the formalization in Mathlib and prevent the agent from reinventing definitions, we passed the agent a Lean file at the start of the process containing relevant definitions for that section, e.g., dihedral groups, roots of unity, or the field extension $\mathbb{Q}(\sqrt{2})$. The agent could reference these definitions and was required to add each formalized statement directly to this file, but was explicitly prohibited from introducing new definitions. The agent generated the top 3 formal statements in Lean for each natural language statement and refined each attempt up to 3 times with feedback from the Lean compiler. We then built the dataset by retaining only those pairs of natural language and formal language statements that corresponded to exercises from the source texts.
@@ B.1.2 Example
This is an example of a theorem statement in the AbstractAlgebra dataset, formalized from Exercise 3 in Chapter 1.2 of Dummit and Foote [dummit2004abstract].
@@
/-- Show that every reflection x = sr^i in D_n has order 2. -/
@@ B.2.1 Dataset Generation
The dataset was generated through an iterative human-in-the-loop process combining automated proof synthesis with expert curation. The initial set of 134 theorems were manually extracted from [nielsen2010quantum]. An automated coding agent (Claude Opus [anthropic2025claudeopus]) first generated formal statements and proof attempts for the theorems, producing both complete proofs and partial derivations. A quantum physics expert then reviewed each statement and proof, identifying gaps, correcting errors, and standardizing operator definitions to ensure that each question was well formed and solvable. The final dataset replaces these proofs with sorry statements.
@@ Appendix C Detected Autoformalization Error
As noted in Section LABEL:subsec:results, 19.7% of Numina’s problems were generated using autoformalization models. While these pipelines enable large-scale dataset construction, they occasionally produce ill-posed theorems that cannot be satisfied in Lean.
@@ Appendix E Tactics
Tactic Ax-Prover DeepSeek Kimina
@@
Table 2: Tactics used by Ax-Prover, DeepSeek, and Kimina. An “X” indicates the model uses the tactic.
@@ Appendix F Case Study: Verifying math in classical cryptographic papers
As a concrete example, we focus on the recent (May 2024) cryptographic paper A New Algorithm for Computing Branch Number of Non-Singular Matrices over Finite Fields from arXiv [mishra2024new]. This work introduces a novel algorithm for computing the branch number – a fundamental metric used to assess the strength of block ciphers such as AES [nist2023aes], PRINCE [borghoff2012prince], and Grøstl [gauravaram2008grostl].
@@
For cryptographers, this makes a practical difference: it enables fast evaluation of candidate matrices when designing new lightweight or high-performance ciphers. The authors demonstrate in Theorem 4 [mishra2024new] that their algorithm achieves significant complexity gains over the naive $O(n^{2}q^{n})$ approach for finite fields of order $q\geq 4$ and square matrices of order $n\geq 4$.
@@ F.4 Correctness, Verification, and Takeaways
The outcome of our work was a formally verified certificate ensuring the correctness of Theorem 1. During the course of constructing this certificate, we identified an error in the original proof. Specifically, the issue arises in Step 2 of the proof:
@@
Our formal verification system flagged these issues because it could not establish the truth of the corresponding statements, revealing logical gaps in the proof. Nevertheless, the authors’ final result remains correct despite the error in the proof.
Appendix G Use Case 2: Quantum Cryptography
In what follows, we expand on Sec. LABEL:sec:QKD by citing the lemma and explaining it in detail. We quote lemma 1 as well as its proof from the supplemental materials in Lo-Chau’s 1999 work in full [lo1998unconditional]:  [^5] We refer to the arXiv version, where the lemma is stated explicitly.
We now restate it with the appropriate context. Let $\rho$ be a density matrix on a $2^{2R}$-dimensional Hilbert space. If its fidelity with the ideal $R$-singlet state satisfies $\langle R\text{~singlets}\mid\rho\mid R\text{~singlets}\rangle>1-\delta$ with $0<\delta\ll 1$, then
@@ Appendix G Use Case 2: Quantum Cryptography
While this is the form reproduced in this work from the Lo–Chau analysis [lo1998unconditional, lo1999unconditional], here we manually replace the statement $\delta\ll 1$ with $\delta\leq 1-\frac{1}{2^{2R}-1}$. To justify this step, we reiterate that the notation $\delta\ll 1$, commonly used in physics, indicates that $\delta$ is smaller than any other relevant quantity in the problem—including the one we construct for our purposes, $1-\frac{1}{2^{2R}-1}$. This step highlights a central challenge in translating physical statements into rigorous mathematics.
We now restate Lo-Chau’s proof. The fidelity condition implies that the largest eigenvalue of $\rho$ is at least $1-\delta$. Since von Neumann entropy is Schur-concave, the maximal entropy under this constraint is achieved by the extremal spectrum $\bigl(1-\delta,\tfrac{\delta}{2^{2R}-1},\ldots,\tfrac{\delta}{2^{2R}-1}\bigr)$, which yields the stated bound.
Our Lean certificate for this result follows this reduction and checks the necessary concavity and normalization facts – producing several reusable and useful lemmas, such as results on majorization and the fact the function $\vec{x}\mapsto\sum_{i}x_{i}\cdot\log x_{i}$ is Schur-convex. This Lean certificate can be made available upon request.
In the Lo–Chau security reduction, the lemma turns “almost-perfect EPR pairs” into a quantitative entropy bound that, together with standard information-theoretic tools, guarantees a limit on an eavesdropper’s knowledge of the final key. Machine-checking this step enables principled composition with other verified components in formal QKD security proofs. Thus, Ax-Prover bridges formal reasoning and quantitative quantum information theory: results such as the Lo–Chau entropy bound no longer have to be taken as a black box, but instead become certified components, ready for use in end-to-end formal verification of QKD security proofs.